The area of a rectangle is (4x^2+4x-3). Find the dimensions of the rectangle using factor theorem
 
 
 

Grade:9

1 Answers

Eshan
5 years ago
Dear student,

We can factorize the given area in the following way-

4x^2+4x-3=4x^2+6x-2x-3=2x(2x+3)-1(2x+3)=(2x-1)(2x+3)
Hence the length and breadth of the rectangle are(2x+3)and(2x-1), respectively.